### v1

@@ Abstract
Early multi-wavelength observations offer crucial insights into the nature of the relativistic jets responsible for gamma-ray bursts and their interaction with the surrounding medium. We present data of GRB 240825A from 17 space- and ground-based telescopes/instruments, covering wavelengths from NIR/optical to X-ray and GeV, and spanning from the prompt emission to the afterglow phase triggered by Swift and Fermi. The early afterglow observations were carried out by SVOM/C-GFT, and spectroscopic observations of the afterglow by GTC, VLT, and TNG determined the redshift of the burst () later. A comprehensive analysis of the prompt emission spectrum observed by Swift-BAT and Fermi-GBM/LAT reveals a rare and significant high-energy cutoff at 76 MeV. Assuming this cutoff is due to absorption allows us to place an upper limit on the initial Lorentz factor, . The optical/NIR and GeV afterglow light curves be described by the standard external shock model, with early-time emission dominated by a reverse shock (RS) and a subsequent transition to forward shock (FS) emission. Our afterglow modelling yields a consistent estimate of the initial Lorentz factor (). Furthermore, the RS-to-FS magnetic field ratio () indicates that the reverse shock region is significantly more magnetized than the FS region. An isotropic-equivalent kinetic energy of erg is derived, and the corresponding -ray radiation efficiency is estimated to be = 3.1%. On the other hand, the standard afterglow model can not reproduce the X-ray light curve of GRB 240825A, calling for improved models to characterize all multi-wavelength data.
@@
Gamma-ray bursts (GRBs) are the most luminous explosions observed in the universe. Over more than half a century of research in this field, the cosmological origins and the demands of extreme relativistic conditions have been acknowledged as fundamentally significant (Ruderman 1975; Metzger et al. 1997). The GRB jet, which is either Poynting-flux-dominated or matter-dominated, undergoes acceleration through the dissipation of its magnetic or thermal energy and reaches relativistic velocities. For prompt emission, the physical origin of the initial emission, i.e. the prompt phase, remains debated, with several scenarios proposed to explain it, such as internal shocks (e.g. Rees & Meszaros 1994; Paczynski & Xu 1994; Kobayashi et al. 1997; Daigne & Mochkovitch 1998), magnetic dissipation (e.g. Spruit et al. 2001; Drenkhahn & Spruit 2002; Giannios 2008; Zhang & Yan 2011; McKinney & Uzdensky 2012; Kumar & Crumley 2015; Sironi et al. 2016; Beniamini & Granot 2016; Granot 2016), and photospheric emission (e.g. Goodman 1986; Thompson 1994; Ghisellini & Celotti 1999; Pe’er et al. 2006; Thompson et al. 2007), etc. The afterglow originates from external dissipation, specifically external shocks sweeping up mass from the circumburst medium, which may include both forward shock (FS) and reverse shock (RS) components (Mészáros & Rees 1997, 1999; Sari & Piran 1999). Therefore, complete observations from prompt emission to afterglow are particularly crucial for understanding the physics of GRBs, including the evolution of their Lorentz factors, the properties of their surrounding environments, and other related aspects. Among them, the relatively rare optical components arising from the reverse shock (RS) can be used as probes to investigate the physics of GRB jets (Fan et al. 2002; Zhang et al. 2003), particularly in constraining the magnetization of the FS and RS regions (Fan et al. 2004; Zhang & Kobayashi 2005).
@@
In this paper, we present GRB 240825A, a GRB with extensive multi-wavelength observations covering both the prompt emission and the afterglow phases. This event was triggered and localized by Swift (Gupta et al. 2024). Follow-up observations of the early and late phases of the afterglow were conducted using Chinese Ground Follow-up Telescope (C-GFT) and the onboard instrument VT (Visible Telescope) of the SVOM mission111https://www.svom.eu/. The SVOM mission, dedicated to the study of gamma-ray bursts, was successfully launched on June 22, 2024, from the Xichang Satellite Launch Center. Notably, SVOM/ECLAIRs did not trigger on this burst; the instrument was not collecting data at the time, as it was inside the South Atlantic Anomaly (SAA) of the Earth’s magnetic field. Furthermore, the burst was located outside ECLAIRs’ field of view and, being in Earth occultation, was also not visible to SVOM/GRM. Additionally, observations from a global network of ground-based telescopes, such as HMT, T80, PAT17, ALT100C, LCOGT, NOT, GTC, VLT, Asiago, REM, TNG, LBT, and MISTRAL@OHP, have contributed to a comprehensive optical photometric and spectral dataset. This multi-facility campaign provides exceptional time-domain and multi-band coverage for in-depth afterglow analysis. In the high-energy prompt emission spectrum of this GRB, a relatively rare high-energy spectral cutoff was detected. According to previous statistical studies (Tang et al. 2015; Li et al. 2023), only 14 such cases of high-energy spectral cutoffs have been reported. The high-energy spectral cutoff provides approximate constraints for estimating the initial Lorentz factor (Krolik & Pier 1991; Fenimore et al. 1993; Woods & Loeb 1995; Baring & Harding 1997; Lithwick & Sari 2001; Baring 2006; Ackermann et al. 2011), offering a valuable comparison to the value derived from afterglow modeling. Moreover, the well-sampled multi-wavelength observations, covering the entire evolution of the afterglow, offer an excellent foundation for deriving the microphysical parameters of this burst.
@@
Section 2 presents a comprehensive overview of the processing and analysis of multi-wavelength observational data, spanning -ray, X-ray, and optical bands. Section 3 details the multi-wavelength data fitting and the inference of microphysical parameters based on the afterglow model, incorporating updated priors from the prompt emission. Using the inferred results, we compare the parameter which describes the magnetic ratio between the forward shock (FS) and reverse shock (RS) regions, as well as the initial Lorentz factor with previous statistical studies. Section 4 provides a summary and discussion of the analysis results for this GRB, covering both the prompt emission and the afterglow phases. Throughout this paper, we adopt the standard cosmological model as km s-1Mpc-1, and (Bennett et al. 2014). And the errors correspond to a 1 confidence level by default.
@@ 2 Observations
GRB 240825A triggered both the Swift Burst Alert Telescope (BAT, Barthelmy et al. 2005) and the Fermi Gamma-Ray Burst Monitor (GBM, Meegan et al. 2009) nearly simultaneously at 15:52:59 UT on 25 August 2024, with BAT being triggered 1 second earlier (Gupta et al. 2024; Fermi GBM Team 2024). Approximately 83 seconds after the trigger, Swift X-Ray Telescope (XRT, Burrows et al. 2005) and Ultra-Violet/Optical Telescope (UVOT, Roming et al. 2005) began follow-up observations. Within the XRT error circle of 12.3 arcseconds, a bright optical afterglow was detected at the coordinates RA = 344.57192 deg, DEC = 1.02686 deg(Gupta et al. 2024). Notably, SVOM/C-GFT commenced follow-up observations only 66 seconds after the trigger (SVOM/C-GFT Team et al. 2024), achieving a response time comparable to that of space-based facilities. As such, this event serves as an exemplary case of successful ground-space coordinated observations and embodies a key vision of the SVOM mission (Wei et al. 2016). The following sections provide a detailed description of the data processing and analysis for the -ray, X-ray, and optical observations.
@@ 2.1.1 Prompt emission phase
During the prompt emission phase, the -ray monitors onboard Swift and Fermi (i.e., BAT and GBM) observed GRB 240825A, and the Large Area Telescope (LAT, Atwood et al. 2009) on Fermi further detected high-energy photons from this source. We conducted a joint analysis of the detected data from these three instruments. For Swift, BAT is a hard X-ray telescope that uses a coded mask to provide a wide field of view and localization, and energy band cover from 15 keV to 350 keV (Barthelmy et al. 2005). The data reduction for BAT followed the standard official analysis threads 222https://www.swift.ac.uk/analysis/bat/index.php to generate the light curve, spectrum and response files. For Fermi, GBM comprises 12 thallium-activated sodium iodide (NaI(Tl)) scintillation detectors and two bismuth germanate (BGO) scintillation detectors (Meegan et al. 2009). Based on the pointing direction of each detector and the position of the source, we selected the data from one NaI detector (n7) and one BGO detector (b1) for the analysis of this event. The Fermi-LAT instrument is a high-performance -ray telescope for the photon energy range from 30 MeV to more than 300 GeV (Atwood et al. 2009). The data reduction for Fermi GBM and LAT was performed using the GBM Data Tools v1.1.1 (Goldstein et al. 2021) and Fermitools v2.2.0 package333https://github.com/fermi-lat/Fermitools-conda/, respectively. The GBM response files were obtained from the Fermi FTP server 444https://heasarc.gsfc.nasa.gov/FTP/fermi/ and are specific to each event, whereas the LAT data were analyzed using the P8R3_TRANSIENT020_V3 instrument response function (IRF).
@@
where and are the evidences from two different models, respectively. Here, the model with the additional power-law component achieved higher evidence, and the Bayesian factor BF showed 83 indicating that the data provides stronger support for this model (Thrane & Talbot 2019b; Jeffreys 1998). The energy peak in spectrum of the Band component is keV, with a low-energy spectral index = and a high-energy spectral index = . The spectral index of the second powerlaw component is = , and the high-energy spectral cutoff in spectrum is MeV.
@@
where is the luminosity distance, ( and are the energy band of the detector) is the correction factor (Bloom et al. 2001), the isotropic -ray energy is . This result is consistent with the report from Konus-Wind (Frederiks et al. 2024) and falls within the range of Type II GRBs (Zhang 2006; Zhang et al. 2009) in the Amati relation (Amati et al. 2002), as shown in Figure 3. In addition, we noticed that the other two papers provided a more detailed analysis and discussion of the prompt emission (Zhang et al. 2025; Wang et al. 2025).
@@ 2.1.2 Extended high-energy observations
Swift-XRT and Fermi-LAT detected emission with extended duration from this source, ranging from X-rays to -rays (Gropp et al. 2024; Di Lalla et al. 2024). We reprocessed the long-duration Swift-XRT data, selecting grades 0–12 for the Photon Counting (PC) mode and grades 0–2 for the Windowed Timing (WT) mode, and restricting the energy range to 0.3–10 keV. The light curves were rebinned using the step-binning method implemented in xrtgrblc. Time-resolved spectral information was taken into account when converting the count rates to energy fluxes, based on the spectral evolution data provided by the online burst analyzer tool666https://www.swift.ac.uk/burst_analyser/ (see more details in Evans et al. 2007, 2009). For the Fermi-LAT data reduction, we selected photon events within the energy range of 100 MeV to 100 GeV and the 0–2200 s interval, using the TRANSIENT event class and the FRONT+BACK type. To minimize contamination from the Earth’s limb, photons with zenith angles exceeding 100∘ were excluded. Subsequently, we identified good time intervals by applying the quality filter condition ((DATA_QUAL1 && LAT_CONFIG==1)). In the standard unbinned likelihood analysis procedure777https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/, we selected a 10∘ region centered on the location of GRB 240825A as the region of interest (ROI). The initial model for the ROI region, generated by the make4FGLxml.py script888http://fermi.gsfc.nasa.gov/ssc/data/analysis/user/, includes the galactic diffuse emission template (gll_iem_v07.fits), the isotropic diffuse spectral model for the TRANSIENT data (iso_P8R3_TRANSIENT020_V3_v1.txt) and all the Fourth Fermi-LAT source catalog (gll_psc_v31.fit; Abdollahi et al. 2020) sources. We defined GRB 240825A as a point source in the model file, with the model set to PowerLaw2999https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/source_models.html, which is simple power law with the integral number of counts between two energies as the normalization. An automatic rebinning algorithm was employed, ensuring at least 10 photons per bin and TS 16 to guarantee that the flux in each bin is significant. The results of the data analysis in this section are used for the multi-band afterglow fitting in Section3.
@@ 2.2.2 SVOM/VT
The VT (Visible telescope) is an optical telescope onboard the Space Variable Objects Monitor mission (SVOM, Wei et al. 2016). The effective aperture is 43 cm. The field of view is 2626 square arcminute, giving the pixel scale of 0.76 arcseconds. It conducts the observation with two channels VT_B and VT_R simultaneously, covering the wavelength of 400-650 nm and 650-1000 nm respectively. Detailed information on VT could be referred to Qiu et al. (2025)(in preparation). During the commission phase, GRB 240825A was observed by VT three epochs via Target Of Opportunity (ToO) mode on August 26, 2024, September 8, 2024 and September 23, 2024. The exposure time was 20 seconds per frame, with a total of 405 frames acquired for VT_B and 434 frames for VT_R band, corresponding to about minutes of observation time. All the data were processed in a standard manner using the IRAF111111IRAF is distributed by the National Optical Astronomical Observatories, which are operated by the Association of Universities for Research in Astronomy, Inc., under cooperative agreement with the National Science Foundation. package., including zero correction, dark correction, and flat-field correction. After the pre-processing, the images for each band obtained during each epoch were stacked to increase the signal-to-noise ratio. The counterpart was clearly detected in all stacked images. Aperture photometry was adopted. The aperture radius and full aperture radius set to be 1.5 pixels and 7.0 pixels, respectively. For each image aperture correction was done to derive the final brightness. The photometric results are listed in Table 3, in which the magnitudes were calibrated in AB magnitude and not corrected for the Galactic extinction.

### v2

@@ Abstract
Early multiwavelength observations offer crucial insights into the nature of the relativistic jets responsible for gamma-ray bursts and their interaction with the surrounding medium. We present data of GRB 240825A from 17 space- and ground-based telescopes/instruments, covering wavelengths from NIR/optical to X-ray and GeV, and spanning from the prompt emission to the afterglow phase triggered by Swift and Fermi. The early afterglow observations were carried out by SVOM/C-GFT, and spectroscopic observations of the afterglow by GTC, VLT, and TNG determined the redshift of the burst () later. A comprehensive analysis of the prompt emission spectrum observed by Swift-BAT and Fermi-GBM/LAT reveals a rare and significant high-energy cutoff at 76 MeV. Assuming this cutoff is due to absorption allows us to place an upper limit on the initial Lorentz factor, . The optical/NIR and GeV afterglow light curves can be described by the standard external shock model, with early-time emission dominated by a reverse shock (RS) and a subsequent transition to forward shock (FS) emission. Our afterglow modeling yields a consistent estimate of the initial Lorentz factor (). Furthermore, the RS-to-FS magnetic field ratio () indicates that the reverse shock region is significantly more magnetized than the FS region. An isotropic-equivalent kinetic energy of erg is derived, and the corresponding -ray radiation efficiency is estimated to be = 3.1%. On the other hand, the standard afterglow model cannot reproduce the X-ray light curve of GRB 240825A, calling for improved models to characterize all multiwavelength data.
@@
Gamma-ray bursts (GRBs) are the most luminous explosions observed in the universe. Over more than half a century of research in this field, the cosmological origins and the demands of extreme relativistic conditions have been acknowledged as fundamentally significant (Ruderman 1975; Metzger et al. 1997). The GRB jet, which is either Poynting-flux-dominated or matter-dominated, undergoes acceleration through the dissipation of its magnetic or thermal energy and reaches relativistic velocities. For prompt emission, the physical origin of the initial emission, i.e. the prompt phase, remains debated, with several scenarios proposed to explain it, such as internal shocks (e.g. Rees & Meszaros 1994; Paczynski & Xu 1994; Kobayashi et al. 1997; Daigne & Mochkovitch 1998), magnetic dissipation (e.g. Spruit et al. 2001; Drenkhahn & Spruit 2002; Giannios 2008; Zhang & Yan 2011; McKinney & Uzdensky 2012; Kumar & Crumley 2015; Sironi et al. 2016; Beniamini & Granot 2016; Granot 2016), and photospheric emission (e.g. Goodman 1986; Thompson 1994; Ghisellini & Celotti 1999; Pe’er et al. 2006; Thompson et al. 2007), etc. The afterglow originates from external dissipation, specifically external shocks sweeping up mass from the circumburst medium, which may include both forward shock (FS) and reverse shock (RS) components (Mészáros & Rees 1997, 1999; Sari & Piran 1999). Therefore, complete observations from prompt emission to afterglow are particularly crucial for understanding the physics of GRBs, including the evolution of their Lorentz factors, the properties of their surrounding environments, and other related aspects. Among them, the relatively rare optical components arising from the RS can be used as probes to investigate the physics of GRB jets (Fan et al. 2002; Zhang et al. 2003), particularly in constraining the magnetization of the FS and RS regions (Fan et al. 2004; Zhang & Kobayashi 2005).
@@
In this paper, we present GRB 240825A, a GRB with extensive multiwavelength observations covering both the prompt emission and the afterglow phases. This event was triggered and localized by Swift (Gupta et al. 2024). Follow-up observations of the early and late phases of the afterglow were conducted using Chinese Ground Follow-up Telescope (C-GFT) and the onboard instrument Visible Telescope (VT) of the SVOM mission111https://www.svom.eu/. The SVOM mission, dedicated to the study of GRBs, was successfully launched on 2024 June 22 from the Xichang Satellite Launch Center. Notably, SVOM/ECLAIRs did not trigger on this burst; the instrument was not collecting data at the time, as it was inside the South Atlantic Anomaly (SAA) of the Earth’s magnetic field. Furthermore, the burst was located outside ECLAIRs’ field of view (FOV) and, being in Earth occultation, was also not visible to SVOM/GRM. Additionally, observations from a global network of ground-based telescopes, such as HMT, T80, PAT17, ALT100C, LCOGT, NOT, GTC, VLT, Asiago, REM, TNG, LBT, and MISTRAL@OHP, have contributed to a comprehensive optical photometric and spectral dataset. This multi-facility campaign provides exceptional time-domain and multiband coverage for in-depth afterglow analysis. In the high-energy prompt emission spectrum of this GRB, a relatively rare high-energy spectral cutoff was detected. According to previous statistical studies (Tang et al. 2015; Li et al. 2023), only 14 such cases of high-energy spectral cutoffs have been reported. The high-energy spectral cutoff provides approximate constraints for estimating the initial Lorentz factor (Krolik & Pier 1991; Fenimore et al. 1993; Woods & Loeb 1995; Baring & Harding 1997; Lithwick & Sari 2001; Baring 2006; Ackermann et al. 2011), offering a valuable comparison to the value derived from afterglow modeling. Moreover, the well-sampled multiwavelength observations, covering the entire evolution of the afterglow, offer an excellent foundation for deriving the microphysical parameters of this burst.
@@
Section 2 presents a comprehensive overview of the processing and analysis of multiwavelength observational data, spanning -ray, X-ray, and optical bands. Section 3 details the multiwavelength data fitting and the inference of microphysical parameters based on the afterglow model, incorporating updated priors from the prompt emission. Using the inferred results, we compare the parameter which describes the magnetic ratio between the FS and RS regions, as well as the initial Lorentz factor with previous statistical studies. Section 4 provides a summary and discussion of the analysis results for this GRB, covering both the prompt emission and the afterglow phases. Throughout this paper, we adopt the standard cosmological model as km s-1Mpc-1, and (Bennett et al. 2014). And the errors correspond to a 1 confidence level by default.
@@ 2 Observations
GRB 240825A triggered both the Swift Burst Alert Telescope (BAT, Barthelmy et al. 2005) and the Fermi Gamma-Ray Burst Monitor (GBM, Meegan et al. 2009) nearly simultaneously at 15:52:59 UT on 2024 25 August, with BAT being triggered 1 second earlier (Gupta et al. 2024; Fermi GBM Team 2024). Approximately 83 seconds after the trigger, Swift X-Ray Telescope (XRT, Burrows et al. 2005) and Ultra-Violet/Optical Telescope (UVOT, Roming et al. 2005) began follow-up observations. Within the XRT error circle of 12.3 arcseconds, a bright optical afterglow was detected at the coordinates RA = 344.57192 deg, DEC = 1.02686 deg(Gupta et al. 2024). Notably, SVOM/C-GFT commenced follow-up observations only 66 s after the trigger (SVOM/C-GFT Team et al. 2024), achieving a response time comparable to that of space-based facilities. As such, this event serves as an exemplary case of successful ground-space coordinated observations and embodies a key vision of the SVOM mission (Wei et al. 2016). The following sections provide a detailed description of the data processing and analysis for the -ray, X-ray, and optical observations.
@@ 2.1.1 Prompt emission phase
During the prompt emission phase, the -ray monitors onboard Swift and Fermi (i.e., BAT and GBM) observed GRB 240825A, and the Large Area Telescope (LAT, Atwood et al. 2009) on Fermi further detected high-energy photons from this source. We conducted a joint analysis of the detected data from these three instruments. For Swift, BAT is a hard X-ray telescope that uses a coded mask to provide a wide field of view and localization, and energy band covers 15 keV to 350 keV (Barthelmy et al. 2005). The data reduction for BAT followed the standard official analysis threads 222https://www.swift.ac.uk/analysis/bat/index.php to generate the light curve, spectrum and response files. For Fermi, GBM comprises 12 thallium-activated sodium iodide (NaI(Tl)) scintillation detectors and two bismuth germanate (BGO) scintillation detectors (Meegan et al. 2009). Based on the pointing direction of each detector and the position of the source, we selected the data from one NaI detector (n7) and one BGO detector (b1) for the analysis of this event. The Fermi-LAT instrument is a high-performance -ray telescope for the photon energy range from 30 MeV to more than 300 GeV (Atwood et al. 2009). The data reduction for Fermi GBM and LAT was performed using the GBM Data Tools v1.1.1 (Goldstein et al. 2021) and Fermitools v2.2.0 package333https://github.com/fermi-lat/Fermitools-conda/, respectively. The GBM response files were obtained from the Fermi FTP server 444https://heasarc.gsfc.nasa.gov/FTP/fermi/ and are specific to each event, whereas the LAT data were analyzed using the P8R3_TRANSIENT020_V3 instrument response function (IRF).
@@
where and are the evidences from two different models, respectively. Here, the model with the additional power-law component achieved higher evidence, and the Bayesian factor BF showed 83 indicating that the data provide stronger support for this model (Thrane & Talbot 2019b; Jeffreys 1998). The energy peak in spectrum of the Band component is keV, with a low-energy spectral index = and a high-energy spectral index = . The spectral index of the second powerlaw component is = , and the high-energy spectral cutoff in spectrum is MeV.
@@
where is the luminosity distance, ( and are the energy bands of the detector) is the correction factor (Bloom et al. 2001), the isotropic -ray energy is . This result is consistent with the report from Konus-Wind (Frederiks et al. 2024) and falls within the range of Type II GRBs (Zhang 2006; Zhang et al. 2009) in the Amati relation (Amati et al. 2002), as shown in Figure 3. In addition, we noticed that the other two papers provided a more detailed analysis and discussion of the prompt emission (Zhang et al. 2025; Wang et al. 2025).
@@ 2.1.2 Extended high-energy observations
Swift-XRT and Fermi-LAT detected emission with extended duration from this source, ranging from X-rays to -rays (Gropp et al. 2024; Di Lalla et al. 2024). We reprocessed the long-duration Swift-XRT data, selecting grades 0–12 for the Photon Counting (PC) mode and grades 0–2 for the Windowed Timing (WT) mode, and restricted the energy range to 0.3–10 keV. The light curves were rebinned using the step-binning method implemented in xrtgrblc. Time-resolved spectral information was taken into account when converting the count rates to energy fluxes, based on the spectral evolution data provided by the online burst analyzer tool666https://www.swift.ac.uk/burst_analyser/ (see more details in Evans et al. 2007, 2009). For the Fermi-LAT data reduction, we selected photon events within the energy range of 100 MeV to 100 GeV and the 0–2200 s interval, using the TRANSIENT event class and the FRONT+BACK type. To minimize contamination from the Earth’s limb, photons with zenith angles exceeding 100∘ were excluded. Subsequently, we identified good time intervals by applying the quality filter condition ((DATA_QUAL1 && LAT_CONFIG==1)). In the standard unbinned likelihood analysis procedure777https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/, we selected a 10∘ region centered on the location of GRB 240825A as the region of interest (ROI). The initial model for the ROI region, generated by the make4FGLxml.py script888http://fermi.gsfc.nasa.gov/ssc/data/analysis/user/, includes the galactic diffuse emission template (gll_iem_v07.fits), the isotropic diffuse spectral model for the TRANSIENT data (iso_P8R3_TRANSIENT020_V3_v1.txt) and all the Fourth Fermi-LAT source catalog (gll_psc_v31.fit; Abdollahi et al. 2020) sources. We defined GRB 240825A as a point source in the model file, with the model set to PowerLaw2999https://fermi.gsfc.nasa.gov/ssc/data/analysis/scitools/source_models.html, which is simple power law with an integral number of counts between two energies as the normalization. An automatic rebinning algorithm was employed, ensuring at least 10 photons per bin and TS 16 to guarantee that the flux in each bin is significant. The results of the data analysis in this section are used for the multiband afterglow fitting in Section3.
@@ 2.2.2 SVOM/VT
The VT (Visible telescope) is an optical telescope onboard the Space-based multiband Variable Objects Monitor (SVOM, Wei et al. 2016) mission. The effective aperture is 43 cm. The FOV is 2626 square arcminute, giving the pixel scale of 0.76 arcseconds. It conducts the observation with two channels VT_B and VT_R simultaneously, covering the wavelength of 400-650 nm and 650-1000 nm respectively. Detailed information on VT could be referred to Qiu et al. (2025)(in preparation). During the commission phase, GRB 240825A was observed by VT three epochs via Target Of Opportunity (ToO) mode on August 26, 2024, September 8, 2024 and September 23, 2024. The exposure time was 20 seconds per frame, with a total of 405 frames acquired for VT_B and 434 frames for VT_R band, corresponding to about minutes of observation time. All the data were processed in a standard manner using the IRAF111111IRAF is distributed by the National Optical Astronomical Observatories, which are operated by the Association of Universities for Research in Astronomy, Inc., under cooperative agreement with the National Science Foundation. package., including zero correction, dark correction, and flat-field correction. After the pre-processing, the images for each band obtained during each epoch were stacked to increase the signal-to-noise ratio. The counterpart was clearly detected in all stacked images. Aperture photometry was adopted. The aperture radius and full aperture radius set to be 1.5 pixels and 7.0 pixels, respectively. For each image aperture correction was done to derive the final brightness. The photometric results are listed in Table 3, in which the magnitudes were calibrated in AB magnitude and not corrected for the Galactic extinction.